# Random Long-Context Access for Mamba via Hardware-aligned Hierarchical Sparse Attention

### v1

@@ 4.2 Long-range Langauge Modeling
We evaluate long-range language modeling on the recognized datasets in the field, which include PG19 [47], ArXiv-math [4], and CodeParrot [62].
@@ Results.
From Table 1, RAMba performs better than the baselines across most datasets, both in-domain and out-domain. In addition, we have two findings. First, Mamba’s perplexity decreases as the context length increases from 4K to 64K, whereas Mamba with attentions show higher perplexity at 64K compared to 16K when not trained with memroy reset. This suggests that Mamba’s memory state may influence the generalization ability of attention mechanisms. Secondly, when trained with memory reset, Mamba with NSA or HSA all exhibit stronger length extrapolation capabilities but show a decline in in-domain performance. This result validates the effectiveness of memory reset, which essentially constrains the model to rely entirely on the content of the text for retrieval, thereby preventing the model from learning shortcuts. Since the model becomes adapted to retrieving information from longer contexts, it is reasonable that its performance declines when handling shorter contexts.
@@ 4.3 Downstream Tasks
We evaluate various models’ long-context modeling abilities on classic tasks like passkey retrieval [34] and the LongBench V2 dataset [7]. To increase task difficulty, we replace numbers in passkey retrieval with random token sequences. Since passkey retrieval is relatively simple, we further fine-tuned the models using synthetic data following RULER [25]. We use a context length of 4K for fine-tuning with a total training step size equivalent to 5% of the pre-training stage. Evaluations were conducted across different lengths on four RULER tasks: Single NIAH (S-N), Multi-queries NIAH (MQ-N), Variable Tracking (VT), and Frequent Words Extraction (FWE). To align with passkey retrieval, keys in Single NIAH were also replaced with random token sequences. We follow Waleffe et al. [59] by using a Cloze format for LongBench evaluation. As LongBench V2 is a zero-shot benchmark thus small models may exhibit randomness, we additionally evaluate on fine-tunable datasets, including summarization like XSUM [38] and CNN [37], and QA tasks like SQuaD [48], HotpotQA [65], and QuALITY [39]. NSA’s implementation currently does not support generative tasks, so its generative results were not reported.
@@ Results.
First, Figure 3 shows that RAMba achieves perfect accuracy in the passkey retrieval task with a 64M context, even without memory reset, while most baselines drop to nearly zero around 64K. However, removing the chunk encoder causes RAMba’s accuracy to decrease as context length grows, highlighting the encoder’s importance for HSA and length generalization. One possible reason is that the hidden states are primarily optimized for predicting the next token, thus requiring an adapter to extract information representing the current context. Second, LongBench’s evaluation results in Table 2 are generally consistent with those of language modeling. However, since LongBench is a zero-shot evaluation dataset, most baselines performed below the random guessing rate of 25%, with the results across different sub-tasks show high variability. Thus, we believe following SFT tasks can provide deeper insights into architectural capabilities for small models.
@@
Table 3 presents the evaluation results of various models fine-tuned on RULER synthetic data across different context lengths. We have four interesting findings. First, HSA demonstrates more precise chunk selection. HSA performs comparably to full-attention on retrieval-related tasks (Single-NIAH, MultiQuery-NIAH, Variable Tracking) over in-domain context length, while NSA lags behind both, supporting our argument that estimating chunk importance using token-to-token attention scores is inaccurate. Meanwhile, HSA achieves strong performance across retrieval tasks despite selecting chunks only once, validating the effectiveness of the hierarchical attention mechanism in learning token-to-chunk relevance. Second, Mamba excels in sequential statistic tasks. While Mamba-based models underperform in retrieval tasks, they outperform in Frequent Word Extraction (FWE), successfully extrapolating up to 256 the pretraining length. This advantage likely stems from Mamba’s continuous memory flow mechanism, which is also partially inherited by RAMba. Third, memory reset helps length generalization. Models with memory reset consistently outperform those without in length extrapolation. Though performance drops significantly beyond 1M context length on more challenging retrieval tasks, still achieving 256 extrapolation. Fourth, Single-NIAH extrapolates only to 4M, while the similar task of passkey retrieval extends to 64M, with the sole difference lying in the prompt pattern as passkey retrieval uses a simpler prompt 222The passkey retrieval template is structured as follows: “(essays) The pass key is ¡PASS KEY¿. (essays) What is the passkey? The passkey is”. The single NIAH template follows this format: “(essays)… One of the special magic numbers for long-context is: ¡PASS KEY¿. (essays)… What is the special magic number for long-context mentioned in the provided text? Answer:” which may facilitate learning more generalizable patterns, underscoring the potential of HSA. Still, precise chunk selection for extremely long contexts remains an open challenge for future work.
@@
Table 4 shows the results of the model after fine-tuning on summarization and generation tasks. RAMba still maintains a lead in the vast majority of tasks, but it are two observations worth noting. First, Mamba with SWA demonstrates an advantage over other models on the SQuaD task. One potential reason is that the SQuaD dataset contains shorter texts even less than the sliding window, making SWA able to randomly access all contexts. This phenomenon suggests that even short-range random access poses certain challenges for RNNs. Second, even though Transformer underperforms Mamba in PPL, its downstream task performance remains strong, second only to RAMba. This further validates the importance of random access capability for downstream tasks.
@@ 4.4 Efficiency Analysis
To evaluate the scalability and efficiency of HSA, we evaluate the runtime and memory footprint of Flash-Attention, NSA, and HSA operators across different sequence lengths, the throughput of various models at different scales, and the per token time consumption during inference. Runtime measurements were conducted with three attention layers only, excluding additional components such as MLPs or Mamba. HSA performed a single chunk selection shared across all three layers, while the sliding window of NSA is still disabled. The memory footprint was reported as a multiple of the Mamba 370M model. We use w/ offloading and w/o offloading to denote whether memory offloading is enabled. To further evaluate the benefits of shared chunk selection, we introduce an ablation group (w/o sharing), where chunk selection and CPU-GPU memory happens in each HSA layer. When measuring training throughput, we enable FSDP [71] and gradient checkpointing [12], running models on 16 Physics Processing Units (PPUs), each with approximately half the computational power of an A100 GPU.
@@ Results.
Figure 4 compares the time consumption and memory footprint of three operators across different context lengths. Both NSA and HSA, as sparse attention mechanisms, significantly outperform Flash-Attention in terms of speed. HSA is faster than NSA because it performs the chunk selection only once, its only operation with quadratic complexity, and shares it across all HSA layers. In contrast, NSA involves computations with quadratic complexity like compressed token attention and chunk selection in every layer, which increases its time consumption. For memory footprint, enabling memory offloading drastically reduces GPU memory usage and slows its growth with increasing context length. Despite introducing CPU-GPU memory exchange, the impact on inference speed is limited, as reported in Table 6. For example, with a chunk size of 64, hidden size of 4096 and batch size of 16, the total offloaded data is , which is acceptable given CPU-GPU bandwidth exceeding 20GB/s. Table 6 shows that RAMba achieves 90% of Mamba’s training throughput. Although HSA is faster than NSA, the additional encoder introduced sometimes offsets this advantage at certain scales. Nonetheless, the results clearly highlight HSA’s high efficiency, excellent scalability during training, and near-constant memory usage during inference.
@@ 5 Conclusion
In this work, we propose a novel hierarchical sparse attention (HSA), which demonstrates excellent training and inference efficiency, while also features advantages such as length extrapolation and long-ragne random access flexibility, making it an ideal architecture for long-text modeling. It achieves perfect accuracy in passkey retrieval with a 64M context, showcasing its potential for extrapolating to infinite lengths. However, in RULER experiments with more complex tasks, accuracy still declines with increasing length, though it significantly outperforms baseline models. How to ensure precise chunk selection on complex tasks over infinite lengths remains a challenge for future research.
Appendix A Hyper-parameters
@@
Appendix B Training hyper-parameters

### v2

@@ Results.
As shown in Table 1, RAMba performs better than the baselines across most datasets, on both in-domain (4K) and out-of-domain (16K & 64K) lengths. In addition, we have two findings. First, Mamba’s perplexity decreases as the context length increases from 4K to 64K, whereas Mamba with attentions shows higher perplexity at 64K compared to 16K when not trained with memory reset. This suggests that Mamba’s memory state may influence the generalization ability of attention mechanisms. Secondly, when trained with memory reset, Mamba with NSA or HSA all exhibit stronger length extrapolation capabilities but show a decline in in-domain performance. This result validates the effectiveness of memory reset, which essentially constrains the model to rely entirely on the content of the text for retrieval, thereby preventing the model from learning shortcuts. Since the model becomes adapted to retrieving information from longer contexts, it is reasonable that its performance declines when handling shorter contexts.
@@ 4.3 Downstream Tasks
We evaluate various models’ long-context modeling abilities on classic tasks like passkey retrieval [35] and the LongBench V2 dataset [7]. To increase task difficulty, we replace numbers in passkey retrieval with random token sequences. Since passkey retrieval is relatively simple, we further fine-tuned the models using synthetic data following RULER [25]. We use a context length of 4K for fine-tuning with a total training step size equivalent to 5% of the pre-training stage. Evaluations were conducted across different lengths on four RULER tasks: Single NIAH (S-N), Multi-queries NIAH (MQ-N), Variable Tracking (VT), and Frequent Words Extraction (FWE). To align with passkey retrieval, keys in Single NIAH were also replaced with random token sequences. We adopt a Cloze format for LongBench evaluation, following Waleffe et al. [62], to address the instruction-following challenges of small models. Since LongBench V2 is a zero-shot benchmark and thus small models may exhibit randomness, we additionally evaluate on fine-tunable datasets, including summarization tasks like XSUM [39] and CNN [38], and QA tasks like SQuaD [49], HotpotQA [68], and QuALITY [40]. NSA’s implementation currently does not support generation, so its results on generative tasks are not reported.
@@ Results.
First, Figure 3 shows that RAMba achieves perfect accuracy in the passkey retrieval task with a 64M context, even without memory reset, while most baselines drop to nearly zero around 64K length. However, removing the chunk encoder causes RAMba’s accuracy to decrease as context length grows, highlighting the encoder’s importance for HSA and length generalization. One possible reason is that the hidden states are primarily optimized for predicting the next token, thus requiring an adapter to extract information representing the current context. Second, LongBench’s evaluation results in Table 2 are generally consistent with those of language modeling. However, since LongBench is a zero-shot benchmark, most baselines performed below the random guess rate of 25%, with the results across different sub-tasks show high variability. Thus, we believe the following SFT tasks can provide deeper insights into architectural capabilities for small models.
@@
Table 3 presents the evaluation results of various models fine-tuned on RULER synthetic data across different context lengths. We have four interesting findings. First, HSA demonstrates more precise chunk selection. HSA performs comparably to full-attention on retrieval-related tasks (S-N, MQ-N, VT) over in-domain context length, while NSA lags behind both, supporting our argument that estimating chunk importance using token-to-token attention scores is inaccurate. Meanwhile, HSA achieves strong performance across retrieval tasks despite selecting chunks only once, validating the effectiveness of the hierarchical attention mechanism in learning token-to-chunk relevance. Second, Mamba excels in sequential statistical tasks. While Mamba-based models underperform in retrieval tasks, they outperform in Frequent Word Extraction (FWE), successfully extrapolating up to 256 the pre-training length. This advantage likely stems from Mamba’s continuous memory flow mechanism, which is also partially inherited by RAMba. Third, memory reset helps length generalization. Models with memory reset consistently outperform those without in length extrapolation. Though performance drops significantly beyond 1M context length on more challenging retrieval tasks, it still achieves 256 extrapolation. Fourth, templates can make a significant difference. Single-NIAH extrapolates only to 4M, while the similar task of passkey retrieval extends to 64M, with the sole difference lying in the template of passkey retrieval being simpler. We elaborate the templates of these two tasks in Appendix D. It suggests that a simpler pattern may facilitate learning more generalizable patterns. Still, precise chunk selection for extremely long contexts remains an open challenge for future work.
@@
Table 4 shows the results of the model after fine-tuning on summarization and generation tasks. RAMba still outperforms in the vast majority of tasks, but two observations are worth noting. First, Mamba with SWA demonstrates advantages over other models on the SQuaD task. One potential reason is that the SQuaD dataset contains shorter texts, even less than the sliding window, making SWA able to randomly access all contexts. This phenomenon shows that integrating random access can improve RNN performance even for short contexts. Second, despite lagging behind Mamba in perplexity (PPL), the Transformer demonstrates strong downstream performance, ranking just behind RAMba. This further validates the importance of random access capability for downstream tasks.
| Models | #params. | LongBench | RULER | |||||||
| Overall | Easy | Hard | Short | Medium | Long | ctx-len=4K | ctx-len=64K | ctx-len=1M | ||
| Mamba-2 | 2.7B | 26.8 | 23.4 | 29.0 | 29.7 | 27.3 | 21.1 | 89.24/26.62/24.68/77.09 | 0.00/0.00/0.00/0.00 | 0.00/0.00/0.00/0.00 |
| RAMba | 2.7B + 110M | 27.3 | 25.1 | 28.6 | 30.4 | 27.3 | 22.1 | 91.00/86.36/96.47/42.95 | 83.58/76.12/100.00/28.36 | 25.00/13.33/50.00/37.14 |
@@ 4.4 Efficiency Analysis
To evaluate the scalability and efficiency of HSA, we measure the runtime and memory footprint of FlashAttention-2, NSA, and HSA operators across different sequence lengths, the throughput of various models at different scales, and the per-token time consumption during inference. Runtime measurements were conducted with three attention layers only, excluding additional components such as MLPs or Mamba. In HSA, a single chunk selection is shared across all three layers, while SWA remains disabled in NSA. The memory footprint is reported as a ratio of the memory occupied by Mamba 370M. We use w/ offloading and w/o offloading to denote whether memory offloading is enabled. To further evaluate the benefits of shared chunk selection, we introduce an ablation group (w/o sharing), where chunk selection and CPU-GPU memory exchanges happen in each HSA layer. When measuring training throughput, we enable FSDP [75] and gradient checkpointing [12], running models on 16 Physics Processing Units (PPUs), each with approximately half the computational power of an A100 GPU.
@@ Results.
Figure 4 compares the time consumption and memory footprint of three operators across different context lengths. Both NSA and HSA, as sparse attention mechanisms, significantly outperform Flash-Attention in terms of speed. HSA is faster than NSA because it performs chunk selection—the only operation with quadratic complexity—only once, and shares it across all HSA layers. In contrast, NSA involves computations with quadratic complexity like compressed token attention and chunk selection in every layer, which increases its time consumption. For memory footprint, enabling memory offloading drastically reduces GPU memory usage and slows its growth with increasing context length. Despite introducing CPU-GPU memory exchange, the impact on inference speed is limited, as reported in Table 7. Table 7 shows that RAMba achieves 90% of Mamba’s training throughput. Although HSA is faster than NSA, the additional encoder sometimes offsets this advantage at certain scales. Nonetheless, the results clearly highlight HSA’s high efficiency, excellent scalability during training, and near-constant memory usage during inference.
@@ 5 Conclusion
In this work, we present Hierarchical Sparse Attention (HSA) and build on it to propose RAMba, which integrates an RNN backbone, length-generalizable sparse attention, and a simple forgetting mechanism. This architecture strikes a strong balance between performance, efficiency, long-range access, and length generalization, offering a foundation for language models with permanent memory.
Appendix A How HSA achieves accurate chunk retrieval
The core principle of previous sparse attentions’ chunk selection lies in approximating the token-to-chunk relevance using unnormalized attention scores. In self-attention, the relevance of token to token is defined as:
where is the dot product between the query of token i and the key of token j.
The relevance of chunk to token is ideally the sum of the relevance of all tokens within that chunk:
where denotes tokens in chunk . However, calculating requires computing (the full softmax normalization across all tokens), which would necessitate full computation and thus undermine the computational efficiency.
To ensure efficiency, they instead approximate chunk relevance using the mean-pooled representation of keys:
where represents the mean-pooling of key representations within the -th chunk. While this approximation bypasses the need for normalized softmax scores across all tokens, it introduces a discrepancy between (the approximation) and (the ideal one).
Let’s consider the following example shown in Figure 5, assuming every 2 tokens form a chunk.
If only the top-2 chunks can be selected, they would choose chunks 2 and 3 according to , thus missing the chunk with the highest sum of attention weights.
Assuming they select the top-2 chunks for sparse attention, the subsequent attention process is shown in Figure 6.
Throughout this process, only participates in chunk selection but not in the forward computation, nor does it receive gradients. The inaccurate chunk selection issue stems from using unnormalized attention logits to estimate chunk importance, with not learnable, making the inaccuracy of chunk selection unavoidable.
For the same example, HSA works as shown in Figure 7
where represents learnable token-to-chunk relevance scores. HSA selects chunks 2 and 3 according to , and applies the hierarchical attention over selected chunks as shown in Figure 8.
In this process, , derived from , participates in the final attention weight computation and the whole forward pass. This allows it to receive gradients and allocate higher weights to chunks with more important tokens. Even if random initialization initially misses the most relevant chunk , continuous training enables the model to gradually learn to select the most relevant chunks.
Appendix B Hyper-parameters
@@
Appendix C Training hyper-parameters
@@
All models are pre-trained on 16 PPUs, with each taking approximately 60 hours.
Appendix D Templates for passkey retrieval and Single-NIAH tasks
The passkey retrieval template is structured as follows: “(essays) The pass key is <PASS KEY>. (essays) What is the passkey? The passkey is”. The single NIAH template follows this format: “(essays)… One of the special magic numbers for long-context is: <PASS KEY>. (essays)… What is the special magic number for long-context mentioned in the provided text? Answer:”
Appendix E RAMba 2.7B
We follow the hyperparameters of Mamba-2 2.7B, where the embedding dimension is 2560, and the total number of layers is 64. We use a two-layer Transformer-based encoder with a hidden size of 2560 and an intermediate dimension of 3392. HSA is inserted into Mamba starting from the 32nd layer, with one HSA layer inserted every 8 layers. In HSA, the group size is 5, with a query head size of 16 for each group. The total additional parameters amount to approximately 110M.
E.1 Post-training
Due to the length generalization issues inherent in the Mamba-2 2.7B, we post-train it to stabilize its perplexity on longer contexts. Specifically, we utilized BPTT for post-tuning the base model. We trained the model on sequences of 32K tokens with a batch size of 16 for 3K steps, totaling 1.5B tokens. This stage takes 5 hours on 32 PPUs.
Then we follow CEPE [69] by freezing most parameters of the Mamba backbone and tuning the parameters of the additional HSA modules for post-training. This process primarily involves the following two stages:
Warmup.
We employ a warmup initialization method by simply training the model to copy the first half of the sentence. Specifically, we append an identical copy of each sentence to itself and train the model to locate the distant context and replicate it. At this stage, all parameters of Mamba-2 are frozen, with only the HSA-related parameters remaining tunable. We train the model with a 32K context length, batch size of 16, for 16K steps, with a peak learning rate of , totaling 8B tokens. This stage takes around 24 hours on 32 PPUs.
Post-Training.
We use LoRA [26] to fine-tune 5% of the parameters in the Mamba-2 module, while keeping all parameters in the HSA module fully trainable. The model is trained with a context length of 32K, a batch size of 16, for 32K steps, using a peak learning rate of , totaling 16B tokens. This stage takes around 48 hours on 32 PPUs.
This model is used for the LongBench evaluation in Table 5.
E.2 RULER finetuning
Since the RULER tasks like NIAH and FWE do not have high requirements for intrinsic knowledge of LLMs, we opt to train RAMba from scratch and fine-tune it on RULER’s synthetic dataset. This approach aims to evaluate whether RAMba trained from scratch can stably converge and demonstrate long-range retrieval capabilities. We conduct pre-training on 60B tokens, which amounts to one-tenth of the Mamba-2 2.7B model, followed by fine-tuning on 1B synthetic data, which takes around 200 hours on 32 PPUs. We also fine-tuned the Mamba-2 2.7B model on the same synthetic dataset for comparison.
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: The claims made in abstract and introduction are supported by experimental results in Section 4.
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: The limitation section is presented in page 16.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification:
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: This paper provides detailed experiment settings in Section 4.1, Appendix A,B, and D, which should be sufficient for reproducing the main experimental results.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [No]
Justification: The sources of all data are explained in the paper, and the code will be open-sourced on GitHub after de-anonymization.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: This paper provides detailed experiment settings in Section 4.1, Appendix A,B, and D, which should be sufficient for reproducing the main experimental results.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [Yes]
Justification: Since each pre-trained model only has a single checkpoint, it is not feasible to evaluate error bars. However, in all experiments, we include multiple datasets or multiple sets of tasks and evaluate using various metrics. Therefore, the statistical significance of the experiments can be verified and supported.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: In Appendix B,D and Section 4.4.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: Our work adheres to all the ethical guidelines outlined by NeurIPS.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [N/A]
Justification:
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification:
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: All dataset details and original authorship are cited in Section 4.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [N/A]
Justification: The code will be published and well documented after de-anonymization.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: Our paper does not involve crowdsourcing nor research with human subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: Our paper does not involve crowdsourcing nor research with human subjects, hence do not require IRB approval.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [N/A]
Justification: Our core method doesn’t involve the usage of LLMs.
Guidelines:
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.